# Emergent Coordination in Multi-Agent Language Models

### v1

@@ Abstract
When are multi-agent LLM systems merely a collection of individual agents versus an integrated collective with higher-order structure? We introduce an information-theoretic framework to test—in a purely data-driven way—whether multi-agent systems show signs of higher-order structure. This information decomposition lets us measure whether dynamical emergence is present in multi-agent LLM systems, localize it, and distinguish spurious temporal coupling from performance-relevant cross-agent synergy. We implement both a practical criterion and an emergence capacity criterion operationalized as partial information decomposition of time-delayed mutual information (TDMI). We apply our framework to experiments using a simple guessing game without direct agent communication and only minimal group-level feedback with three randomized interventions. Groups in the control condition exhibit strong temporal synergy but only little coordinated alignment across agents. Assigning a persona to each agent introduces stable identity-linked differentiation. Combining personas with an instruction to “think about what other agents might do” shows identity-linked differentiation and goal-directed complementarity across agents. Taken together, our framework establishes that multi-agent LLM systems can be steered with prompt design from mere aggregates to higher-order collectives. Our results are robust across emergence measures and entropy estimators, and not explained by coordination-free baselines or temporal dynamics alone. Without attributing human-like cognition to the agents, the patterns of interaction we observe mirror well-established principles of collective intelligence in human groups: effective performance requires both alignment on shared objectives and complementary contributions across members.
@@ 1 Introduction
Recent advances in generative AI (specifically LLMs) have led to tremendous advances in multi-agent systems (Qu et al., 2024; Hong et al., 2024; Qian et al., 2023; Li et al., 2024a; Subramaniam et al., 2025). Multi-agent systems often show impressive performance increases over single-agent solutions (Wu et al., 2023; Chen et al., 2023; Li et al., 2024b; Tao et al., 2024). A key argument behind such performance gains are claims to “greater-than-the-sum-of-its-parts” effects from differentiated agents (Chen et al., 2023, p.1). Connecting multiple differentiated agents has the potential benefit to leverage unique contributions that would not be available if the task was assigned to only a single agent (Fazelpour & De-Arteaga, 2022; Tollefsen et al., 2013; Luppi et al., 2024). Conceptually, any group gains depend on emergence and synergy (Theiner, 2018; Riedl et al., 2021; Page, 2008). [^1] This is true for both human groups and multi-agent systems. Despite impressive performance of many multi-agent systems we do not yet have a principled understanding when and how such synergy emerges, what role agent differentiation plays, and how to steer it systematically. This all points to a crucial need for deeper understanding of collective intelligence in multi-agent systems.
In this paper, we address the question when is a population of LLM agents a mere collection vs. a higher-order collective? We follow a purely data-driven approach to assess whether multi-agent systems show signs of emergence in the form of higher-order synergy characterized by structural coupling and joint information about future states and task outcomes. Intuitively, synergy refers to information about a target that a collection of variables provide only jointly but not individually (Rosas et al., 2020; Humphreys, 1997). We develop a principled framework (Figure 1a) building on new insights in information theory (Rosas et al., 2020; Mediano et al., 2022). Our goal is to develop methods to measure emergent role specialization, determine whether multi-agent systems show such signs of emergent synergy, explore whether synergy enables increased performance, and explore whether emergent behavior can be systematically steered with prompting.
We apply our framework to study multi-agent systems of gpt-4.1 and Llama-3.1-8B agents solving a simple group guessing task (Figure 1b) with three treatment interventions: a control condition, a condition that assigns personas to each agent, and one that uses personas with an instruction to “think about what other agents might do” (a theory of mind (ToM) prompt). The paper is organized into addressing three concrete research questions:
@@ 3. RQ3:
    Can we design prompts, roles, or reasoning structures that steer the internal coordination of multi-agent to encourage positive, goal-directed synergy?
Our findings provide evidence for multi-agent LLM system capacity for emergence and that it underpins performance. Through a variety of complementary analyses and robustness tests, we show that coordination style across interventions is very different. While emergence is present in all, only the ToM-prompt condition leads to groups with identity-linked differentiation and goal-directed complementarity across agents: they operate as an integrated, goal-directed unit. This is consistent with research on human groups showing that performance gains are not automatic (e.g., Stasser & Titus, 1985) and that complementarity needs to be balanced with integration and goal alignment (Theiner, 2018; DeChurch & Mesmer-Magnus, 2010; Riedl et al., 2021).
This paper helps us understand when and how multi-agent systems exhibit higher-order properties, how to control them, and their internal coordination. This can inform multi-agent system design by showing how to combine agents effectively. We make four contributions:
  1. 1.
@@
  3. 3.
    Demonstrate how to steer emergence with specific prompts. Prompt-level manipulations causally change higher-order dependencies and reliably induces distinct coordination regimes, shifting collectives from spurious and misdirected synergy to stable & goal-aligned complementarity driven by differentiated identities.
  4. 4.
@@ 2 Method
We study a group guessing game (developed by Goldstone et al., 2024, who called it group binary search) without communication: agents propose integers whose sum needs to match a randomly generated hidden target number. Agents are unaware of each others’ guesses and the size of their group, and receive only group-level feedback “too high” or “too low.” The task is challenging because identical strategies induce oscillation and only complementary strategies yield success. This setting naturally pits redundancy (alignment) against synergy (useful diversity), and admits clear nulls via row‑wise (break identities) and column‑wise (break alignment) surrogates. Research on humans has shown that groups can reliably solve this task through the emergence of specialized, complementary roles (Goldstone et al., 2024).
@@
How would we know if a multi-agent system shows emergent properties that could suggest that the sum is more than its parts? We develop a principled test based on a recent formal theory of dynamic emergence based on information decomposition (Rosas et al., 2020; Mediano et al., 2022; Bedau & Humphreys, 2008). Our framework connects emergence with information about a system’s temporal evolution—future states of the whole—with information that cannot be traced to the current state of its parts. This framework is based on partial information decomposition (PID; Williams & Beer, 2010) and time-delayed mutual information (TDMI; Luppi et al., 2022), is data-driven, quantifiable, and amenable to empirical testing with falsifiable conjectures against specific null hypotheses. Combining this with permutation tests, focused either on breaking connections between agents or connections across time allow us not only to quantity emergent properties but also to localize them in the system. This is crucial for multi-agent systems because we care about distinguishing synergy among specialized, differentiated agents from mere dynamic alignment.
We implement three tests. The first, termed the practical criterion, concerns predicting a macro signal $V$ of the system. It asks whether the macro contains predictive information beyond any individual part. Given microstate $X_{t}=(X_{1,t},\dots,X_{n,t})$ and macro $V_{t}=f(X_{t})$. We align samples $(t,t+\ell)$ and compute the score as
$S_{\mathrm{macro}}(\ell)=I(V_{t};V_{t+\ell})-\sum_{k=1}^{n}I(X_{k,t};V_{t+\ell}).$
@@
The second test, termed emergence capacity, captures ability of the multi-agent system to host any synergy. It measures the predictive synergy from two agents’ current states to their joint future state. For each pair of agents $(i,j)$ and time $t$, let the sources be $X_{i,t}$ and $X_{j,t}$, and define the bivariate target as the next-step joint state $T_{ij,t+\ell}\equiv(X_{i,t+\ell},X_{j,t+\ell})$. From the joint contingency table over $(X_{i,t},X_{j,t},T_{ij,t+\ell})$ we compute a two-source PID of the predictive information
$I(\{X_{i,t},X_{j,t}\};\,T_{ij,t+\ell})=\mathrm{UI}_{i}+\mathrm{UI}_{j}+\mathrm{Red}_{ij}+\mathrm{Syn}_{ij}$
and take $\mathrm{Syn}_{ij}$ as the pairwise dynamical synergy. A positive score $\mathrm{Syn}_{ij}>0$ indicates that the macro captures predictive information about the joint future not recoverable from any single component. We compute this for all unordered pairs $(i,j)$ and take the median as group-level synergy capacity. Compared to the practical criterion, this test is limited to detect synergy of order $k=2$ but has the advantage that it does not rely on the definition of a suitable whole-system macro signal.
Finally, we implement a coalition test. Let $I_{3}$ be the mutual information between a triplet’s current state and the future macro signal: $I_{3}=I((X_{i,t},X_{j,t},X_{k,t});V_{t+\ell})$. I.e., how much the three agents jointly predict the macro’ future. High $I_{3}$ indicates that agents’ behaviors are coherently organized toward the shared macro, whereas low $I_{3}$ signals weak alignment or uncoordinated behavior. Then
$G_{3}=I_{3}-\textit{max}(I_{2\{1,2\}},I_{2\{1,3\}},I_{2\{2,3\}})$
measures the additional predictive information the full triplet provides over the most predictive pair $I_{2}$ (a form of whole-minus-parts metric; Barrett & Seth, 2011; Mediano et al., 2025). If we find $G_{3}>0$ this means no pair is sufficient to capture information that the triplet contains about the macro signal at $t+\ell$. We again compute $I_{3}$ and $G_{3}$ for all possible triplets and take the median as group measure. This measure complements the other two by offering a coalition-level test geared towards functional relevance: it helps answers whether the joint information provided by coalitions of agents is actually about the goal. It allows us to localize where macro predictability depends on beyond-pair structure and to rule out “explained by best pair” explanations.
@@
We use the Williams–Beer two‑source PID (Williams & Beer, 2010) with the $I_{min}$ (minimum specific information) redundancy, estimated via plug‑in probabilities. For robustness, we also use Jeffreys-smoothed probabilities, and alternatively MMI redundancy (minimum of whole mutual informations; Mediano et al., 2025) with MI estimated with Miller–Madow bias correction (Miller, 1955). All variables are discretized into $K=2$ bins using quantile binning. Data are encoded as factors with fixed levels 1,2{1,2} (and for the joint target as the Cartesian product levels) to avoid dropping empty categories. We report $\ell=1$ which is suitable to detect next-step oscillating behavior (see task description). Because trials end at success, episode lengths vary (see plot of rounds to success in Appendix). We report additional sensitivity tests using “early synergy” truncated to a fixed horizon of $H\in\{10,15\}$ rounds to ensure comparability (results in Appendix A.2 and A.11; main analyses use full trajectories).
@@
Small-sample entropy estimation is often challenging because the true distribution is only partially observed and many outcome categories receive few or even zero counts (“empty bins”). Such finite-sample estimation of information measures is biased upward and can yield spurious positive findings. Bias grows with more bins, dimensionality (e.g., using triplets instead of pairs), and small $N$ (few timesteps). The empirical plug-in estimator replace true probabilities with sample frequencies and unseen (or under-sampled) outcomes often are assigned zero or very low probabilities (Hausser & Strimmer, 2009). To account for this issue, we follow best practices and take several steps. First, we compute the emergence capacity and the coalition tests as order $k=2$ instead of the entire system of $n$ agents. Second, we apply quantile binning (with two bins) to reduce the dimensionality of our data (sensitivity analyses with three bins reported in A.8). Third, we use bias-corrected entropy estimator with Jeffreys’ prior, which smooths estimates and avoids empty bins by adding $\alpha=\frac{1}{2}$ pseudo-counts in the Dirichlet estimator (Jeffreys, 1946). This reduces systematic inflation and cures zero‑count pathologies. Finally, as a robustness check, we also report results using the Miller–Madow bias-corrected estimator (Miller, 1955), and the MMI redundancy measure (MMI typically overestimates redundancy, making synergy estimates more conservative; Mediano et al., 2025).
@@
We complement the information theoretic measures with a test for agent differentiation based on hierarchical (mixed) models (Gelman & Hill, 2007). We estimate a sequence of three models for each multi-agent experiment:
$m_{0}:$ $\quad y_{i}=\beta_{0}+u_{\text{time}[i]}+\epsilon_{i}$
$m_{1}:$ $\quad y_{i}=\beta_{0}+u_{\text{time}[i]}+u_{\text{agent}[i]}+\epsilon_{i}$
$m_{2}:$ $\quad y_{i}=\beta_{0}+u_{\text{time}[i]}+u_{\text{agent}[i],0}+u_{\text{agent}[i],\text{time}[i]}+\epsilon_{i}$
@@ 3.1 Preliminary Experiments: Group Size and Temperature
To see if LLMs can solve this task and how sensitive results are to different group sizes and temperature settings we ran a first set of preliminary experiments (see Appendix for prompt). We used a frontier-class model, OpenAI’s gpt-4.1-2025-04-14. We varied group size from 3 to 15, and temperature from $[0,1]$ in steps of $0.1$. For each grid point we ran 50 group experiments (13 group sizes $\times$ 11 temperature settings $\times$ 50 groups = 7,150 experiments; Figure 1c). Multi-agent systems of gpt-4.1 agents can solve the task reliably but encounter substantial challenges. We fit a logistic regression, predicting the probability of success from group size and temperature. We find the task is significantly easier for smaller groups: each additional group member decreased the odds by roughly 8% (OR = 0.92, $p<10^{-16}$). This replicates the same results as for human groups (Goldstone et al., 2024). Each unit increase in temperature increased the odds of success by approximately 50% (OR = 1.50, $p<10^{-7}$).

### v2

@@ Abstract
When are multi-agent LLM systems merely a collection of individual agents versus an integrated collective with higher-order structure? We introduce an information-theoretic framework to test—in a purely data-driven way—whether multi-agent systems show signs of higher-order structure. This information decomposition lets us measure whether dynamical emergence is present in multi-agent LLM systems, localize it, and distinguish spurious temporal coupling from performance-relevant cross-agent synergy. We implement a practical criterion and an emergence capacity criterion operationalized as partial information decomposition of time-delayed mutual information (TDMI). We apply our framework to experiments using a simple guessing game without direct agent communication and minimal group-level feedback with three randomized interventions. Groups in the control condition exhibit strong temporal synergy but little coordinated alignment across agents. Assigning a persona to each agent introduces stable identity-linked differentiation. Combining personas with an instruction to “think about what other agents might do” shows identity-linked differentiation and goal-directed complementarity across agents. Taken together, our framework establishes that multi-agent LLM systems can be steered with prompt design from mere aggregates to higher-order collectives. Our results are robust across emergence measures and entropy estimators, and not explained by coordination-free baselines or temporal dynamics alone. Without attributing human-like cognition to the agents, the patterns of interaction we observe mirror well-established principles of collective intelligence in human groups: effective performance requires both alignment on shared objectives and complementary contributions across members.
@@ 1 Introduction
Recent advances in generative AI (specifically LLMs) have led to tremendous advances in multi-agent systems (Qu et al., 2024; Hong et al., 2024; Qian et al., 2023; Li et al., 2024a; Subramaniam et al., 2025). Multi-agent systems often show impressive performance increases over single-agent solutions (Wu et al., 2023; Chen et al., 2023; Li et al., 2024b; Tao et al., 2024). A key argument behind such performance gains are claims to “greater-than-the-sum-of-its-parts” effects from differentiated agents (Chen et al., 2023, p.1). Connecting multiple differentiated agents has the potential benefit to leverage unique contributions that would not be available if the task was assigned to a single agent (Fazelpour and De-Arteaga, 2022; Tollefsen et al., 2013; Luppi et al., 2024; Lix et al., 2022). Conceptually, any group gains depend on synergy and emergence (Theiner, 2018; Riedl et al., 2021; Page, 2008; Hayek, 1945). [^1] This is true for both human groups and multi-agent systems. Despite impressive performance of many multi-agent systems we do not yet have a principled understanding when and how such synergy arises, what role agent differentiation plays, and how to steer it systematically. This all points to a crucial need for deeper understanding of collective intelligence in multi-agent systems, specifically whether multi-agent LLMs exhibit any synergy or complementarity at all—the necessary prerequisite for any absolute team-over-solo gains.
Determining whether multi-agent systems function as genuine collectives—rather than merely aggregating agents—requires a principled measure of synergy. We follow a purely data-driven approach to assess whether these systems exhibit higher-order synergy, characterized by structural coupling and joint information about future states and task outcomes, as evidence of emergent collective behavior. Intuitively, synergy refers to information about a target that a collection of variables provide only jointly but not individually (Rosas et al., 2020; Humphreys, 1997). [^2] The key claim explored in this paper is conditional, cross-agent synergy—i.e., synergistic information and coordinated differentiation across agents given the multi-agent constraint (not absolute outperformance over a solo agent). We do not attempt to establish team-over-solo superiority on this task. We develop a principled framework (Figure 1a) building on new insights in information theory (Rosas et al., 2020; Mediano et al., 2022b). Our goal is to develop methods to measure emergent role specialization, determine whether multi-agent systems show such signs of emergent synergy, explore whether synergy enables increased performance, and explore whether emergent behavior can be systematically steered with prompting. [^3] See GitHub for replication code https://github.com/riedlc/AI-GBS.
We apply our framework to study multi-agent systems of GPT-4.1, Llama-3.1-8B, Llama-3.1-70B, Gemini 2.0 flash, and Qwen3 agents solving a simple group guessing task (Figure 1b) with three treatment interventions: a control condition, a condition that assigns a persona to each agent, and one that uses personas with an instruction to “think about what other agents might do” (a theory of mind (ToM) prompt). The paper is organized into addressing three concrete research questions:
@@ 3. RQ3:
    Can we design prompts, roles, and reasoning structures that steer the internal coordination regimes of multi-agent to encourage outcome-relevant, goal-directed synergy?
Our findings provide evidence for multi-agent LLM system capacity for emergence and that it underpins performance. Through a variety of complementary analyses and robustness tests, we show that coordination style across interventions is very different. While emergence is present in all, only the ToM-prompt condition leads to groups with identity-linked differentiation and goal-directed complementarity across agents: they operate as a dynamically stable, integrated, goal-directed unit. This is consistent with research on human groups showing that performance gains are not automatic (e.g., Stasser and Titus, 1985) and that complementarity needs to be balanced with integration and goal alignment (Theiner, 2018; DeChurch and Mesmer-Magnus, 2010; Riedl et al., 2021).
This paper helps us understand when and how multi-agent systems exhibit higher-order properties, their internal coordination structure, and how to control them. This can inform multi-agent system design by showing how to combine agents effectively. We make four contributions:
@@
  • Demonstrate how to steer emergence with specific prompts. Prompt-level manipulations causally change higher-order dependencies and reliably induces distinct coordination regimes, shifting collectives from spurious and misdirected synergy to stable and goal-aligned complementarity driven by differentiated identities.
@@ 2 Method
We study a group guessing game (developed by Goldstone et al. (2024) who called it “group binary search”) without communication: agents propose integers whose sum needs to match a randomly generated hidden target number. Agents are unaware of each others’ guesses and the size of their group, and receive only group-level feedback “too high” or “too low.” The task is challenging because identical strategies induce oscillation and only complementary strategies yield success. This setting naturally pits redundancy (alignment) against synergy (useful diversity), and admits clear nulls via row‑wise (break identities) and column‑wise (break alignment) surrogates. The binary search setting is a minimalist testbed to isolate emergence of cross-agent complementarity and dynamic coordination, not to establish team-over-solo superiority. The implementation of a textbook-like XOR dynamic allows us to study the benefits of emergence capacity conditional on a multi-agent setup (and not to establish team-over-solo superiority). Research on humans has shown that groups can reliably solve this task through the emergence of specialized, complementary roles (Goldstone et al., 2024).
@@
How would we know if a multi-agent system shows emergent properties that could suggest that the sum is more than its parts? We develop a principled test based on a recent formal theory of dynamic emergence based on information decomposition (Rosas et al., 2020; Mediano et al., 2022b; Bedau and Humphreys, 2008). Our framework connects emergence with information about a system’s temporal evolution—future states of the whole—with information that cannot be traced to the current state of its parts. This framework is based on partial information decomposition (PID; Williams and Beer, 2010) and time-delayed mutual information (TDMI; Luppi et al., 2022), is data-driven, quantifiable, and amenable to empirical testing with falsifiable conjectures against specific null hypotheses. Combining this with permutation tests, focused either on breaking connections between agents or connections across time allow us not only to quantity emergent properties but also to localize them in the system. This is crucial for multi-agent systems because we care about distinguishing synergy among specialized, differentiated agents from mere dynamic alignment.
We implement three tests. The first, termed emergence capacity, captures ability of the multi-agent system to host any synergy. It measures the predictive synergy from two agents’ current states to their joint future state, quantified using mutual information $I(\cdot;\cdot)$. For each pair of agents $(i,j)$ and time $t$, let the sources be $X_{i,t}$ and $X_{j,t}$, and define the bivariate target as the joint future state $t+\ell$, $T_{ij,t+\ell}\equiv(X_{i,t+\ell},X_{j,t+\ell})$, where $\ell$ is the integration timescale (Mediano et al., 2022a). From the joint contingency table over $(X_{i,t},X_{j,t},T_{ij,t+\ell})$ we compute a two-source PID of the predictive information
$I(\{X_{i,t},X_{j,t}\};\,T_{ij,t+\ell})=\mathrm{UI}_{i}+\mathrm{UI}_{j}+\mathrm{Red}_{ij}+\mathrm{Syn}_{ij}$ (1)
and take $\mathrm{Syn}_{ij}$ as the pairwise dynamical synergy. A positive score $\mathrm{Syn}_{ij}>0$ indicates that the macro captures predictive information about the joint future not recoverable from any single component. We compute this for all unordered pairs $(i,j)$ and take the median as group-level synergy capacity. This test is limited to detect synergy of order $k=2$ but has the advantage that it does not rely on the definition of a suitable whole-system macro signal.
The second test, termed the practical criterion, concerns predicting a macro signal $V$ of the system. It asks whether the macro contains predictive information beyond any individual part. Given microstate $X_{t}=(X_{1,t},\dots,X_{n,t})$, macro $V_{t}=f(X_{t})$. We align samples $(t,t+\ell)$ and compute the score as
$S_{\mathrm{macro}}(\ell)=I(V_{t};V_{t+\ell})-\sum_{k=1}^{n}I(X_{k,t};V_{t+\ell}).$ (2)
@@
To explore system dynamics, we implement a coalition test, an extension of the practical criterion to triplets. Let $I_{3}$ be the mutual information between a triplet’s current state and the future macro signal: $I_{3}=I((X_{i,t},X_{j,t},X_{k,t});V_{t+\ell})$. I.e., how much the three agents jointly predict the macro’ future serves as a measure of the system’s information-processing capacity (Mediano et al., 2022a). High $I_{3}$ indicates that agents’ behaviors are tightly coupled and contain rich information about the emerging collective outcome, a pattern frequently associated with collective intelligence (De Vincenzo et al., 2017; Wicks et al., 2007; Mora and Bialek, 2011; Barnett et al., 2013). Low $I_{3}$ signals weak alignment, resulting in uncoordinated or chaotic collective behavior. Then
$G_{3}=I_{3}-\textit{max}(I_{2\{1,2\}},I_{2\{1,3\}},I_{2\{2,3\}})$ (3)
measures the additional predictive information the full triplet provides over the most predictive pair $I_{2}$ (a form of whole-minus-parts metric; Barrett and Seth, 2011; Mediano et al., 2025). If we find $G_{3}>0$ this means no pair is sufficient to capture information that the triplet contains about the macro signal at $t+\ell$. We again compute $I_{3}$ and $G_{3}$ for all possible triplets and take the median as group measure. This measure complements the other two by offering a coalition-level test geared towards functional relevance: it helps answers whether the joint information provided by coalitions of agents is actually about the goal. It allows us to localize where macro predictability depends on beyond-pair structure and to rule out “explained by best pair” explanations.
@@
We use the Williams–Beer two‑source PID (Williams and Beer, 2010) with the $I_{min}$ (minimum specific information) redundancy, estimated via plug‑in probabilities. For robustness, we also use Jeffreys-smoothed probabilities, and alternatively MMI redundancy (minimum of whole mutual informations; Mediano et al., 2025) with mutual information estimated with Miller–Madow bias correction (Miller, 1955). All variables are discretized into $K=2$ bins using quantile binning. Data are encoded as factors with fixed levels 1,2{1,2} (and for the joint target as the Cartesian product levels) to avoid dropping empty categories. We report $\ell=1$ which is suitable to detect next-step oscillating behavior (see task description). Because trials end at success, episode lengths vary (see plot of rounds to success in Appendix). We report additional sensitivity tests using “early synergy” truncated to a fixed horizon of $H\in\{10,15\}$ rounds to ensure comparability (results in Appendix A.2 and A.12; main analyses use full trajectories).
@@
Small-sample entropy estimation is often challenging because the true distribution is only partially observed and many outcome categories receive few or even zero counts (“empty bins”). Such finite-sample estimation of information measures is biased upward and can yield spurious positive findings. Bias grows with more bins, dimensionality (e.g., using triplets instead of pairs), and small $N$ (few timesteps). The empirical plug-in estimator replace true probabilities with sample frequencies and unseen (or under-sampled) outcomes often are assigned zero or very low probabilities (Hausser and Strimmer, 2009). To account for this issue, we follow best practices and take several steps. First, we compute the emergence capacity and the coalition tests as order $k=2$ instead of the entire system of $n$ agents. Second, we apply quantile binning (with two bins) to reduce the dimensionality of our data (sensitivity analyses with three bins reported in A.8). Third, we use bias-corrected entropy estimator with Jeffreys’ prior, which smooths estimates and avoids empty bins by adding $\alpha=\frac{1}{2}$ pseudo-counts in the Dirichlet estimator (Jeffreys, 1946). This reduces systematic inflation and cures zero‑count pathologies. Finally, as a robustness check, we also report results using the Miller–Madow bias-corrected estimator (Miller, 1955), and the MMI redundancy measure (MMI typically overestimates redundancy, making synergy estimates more conservative; Mediano et al., 2025).
@@
Our third pillar evaluates agent differentiation via hierarchical (mixed) modeling (Gelman and Hill, 2007), asking whether agents adopt consistent, person-specific patterns. We estimate a sequence of three models for each multi-agent experiment:
$m_{0}\colon$ $\quad y_{i}=\beta_{0}+u_{\text{time}[i]}+\epsilon_{i}$
$m_{1}\colon$ $\quad y_{i}=\beta_{0}+u_{\text{time}[i]}+u_{\text{agent}[i]}+\epsilon_{i}$
$m_{2}\colon$ $\quad y_{i}=\beta_{0}+u_{\text{time}[i]}+u_{\text{agent}[i],0}+u_{\text{agent}[i],\text{time}[i]}+\epsilon_{i}$
@@ 3.1 Preliminary Experiments: Group Size and Temperature
To see if LLMs can solve this task and how sensitive results are to different group sizes and temperature settings we ran a first set of preliminary experiments (see Appendix for prompt). We used OpenAI’s gpt-4.1-2025-04-14. We varied group size from 3 to 15, and temperature from $[0,1]$ in steps of $0.1$. For each grid point we ran 50 group experiments (13 group sizes $\times$ 11 temperature settings $\times$ 50 groups = 7,150 experiments; Figure 1c). Multi-agent systems of GPT 4.1 agents can solve the task reliably but encounter substantial challenges. We fit a logistic regression, predicting success from group size and temperature. We find the task is significantly easier for smaller groups: each additional group member decreased the odds by roughly 8% (OR = 0.92, $p<10^{-16}$), consistent with results as for human groups (Goldstone et al., 2024). Each unit increase in temperature increased the odds of success by about 50% (OR = 1.50, $p<10^{-7}$).